# Lattice Boltzmann methods for soft flowing matter

### v1

@@ V.6 Peta-scale simulations of deep-sea sponges
We conclude this section with a discussion on a recent application where the LB method is adopted to study the fluid dynamics of the deep-sea glass sponge Euplectella aspergillum Falcucci et al. (2021), a soft living organism that, besides representing a fascinating example of life under extreme conditions (at depth of $100$-$1000$ m in the Pacific Ocean and Antarctic area with no ambient sunlight), displays intriguing structural properties due to the trabecular arrangement of its skeletal system Monn et al. (2015); Fernandes et al. (2020).
@@
The interest of this study lies in the fact that, despite the use of a single-component interface-free lattice Boltzmann (thus its easiest realization), the detailed three-dimensional reconstruction of the skeleton of the sponge and the computation of the fluid flow in which this organism is immersed need to leverage two of the main advantages of the method, namely the excellent scalability on parallel architectures and the capability of handling complex geometries. Indeed, to reproduce the sponge’s living conditions, the simulations feature over fifty billion grid points spanning approximately four spatial decades, while a complex mesh structure is used to simulate the skeleton.
@@
In Fig.21, we show a simulation of the sponge immersed in a hydrodynamic flow at $Re\simeq 2000$, where $Re=uD/\nu$, being $u$ the water velocity, $\nu$ is its viscosity and $D$ is the sponge diameter at the top section. The sponge structure comprises an anchoring bulb, a section connecting the bulb to the body, the main body and the final sieve plate located at the top. The body, in particular, is simulated using two intersecting patterns of lattices, the first composed of filaments orthogonally crossing each other and the second made by smaller ligaments placed at $45^{\circ}$ with respect to the first one. The resulting cylindrical structure finally bends into a cone connecting with the anchoring bulb. Also, no-slip conditions are imposed at the internal boundaries of the sponge and at the seafloor.
@@
The simulations show, for example, that a considerable decrease in the flow speed occurs inside the sponge’s body (where vortices are formed) and downstream of the organisms, while intermittent fluid patterns emerge several diameters further away. These features are supposed to play a role in feeding through a selection mechanism where nutrients are absorbed in the sponge chambers and inorganic particles are discarded. Also, at $Re>100$ intermittency patterns are delayed, an effect that reduces the stress and potentially enhances the mechanical stability of the sponge. It has also been noted that the presence of the sponge’s tissue (neglected in these simulations) could impact the structure of the fluid flow because the tissue may form a barrier with low permeability over the sponge wall Leys et al. (2022). These conditions could be partially reproduced by modeling the sponge as a porous medium with an effective resistance to the flow, although additional experiments and empirical data would be needed to test more complex models.
Figure 21: Simulation of a deep-sea sponge immersed in a hydrodynamic flow at $Re\simeq 2000$, where contours of helicity ${\cal H}$ (defined as ${\cal H}={\bf u}\cdot{\bm{\omega}}$, where ${\bm{\omega}}$ is the vorticity) and streaklines of the flow are shown. The figure is reproduced with permission from Ref.Falcucci et al. (2021).
V.7 From applications to engineering design
@@
The set of selected applications presented above is meant to convey the idea of the flexibility and computational efficiency of the LB methods for soft-flowing matter across different regimes of motion. Such flexibility and efficiency are expected to find profitable use in various sectors of industrial and materials design, such as food science, chemical-pharmaceutical (drug design and delivery), tissue engineering and many others. While the details of each and every specific application must necessarily be worked out on a case-by-case basis, in the following we wish to convey the flavor of the potential of computational design in the above sectors based on the LB methods discussed in this review.
@@
For the sake of concreteness, let us refer to a soft material, say a dense emulsion flowing in a box device of height $H=0.1$, width $W=1$ and length $L=10$, all in millimeters. With a lattice spacing $\Delta x=10^{-3}$ (1 micron), the LB simulation consists of $G=10^{2+3+4}=10^{9}$ lattice cells. Such volume contains about a thousand cylindrical droplets of diameter $D=0.1$ mm (100 microns) and height $H$. At a processing speed of $100$ GLUPS, i.e., a hundred billion lattice sites per second, the LB code updates the full box domain a hundred times in a single second of computer time With a timestep of $\Delta t=10^{-9}$ (1 ns), one million timesteps cover $1$ millisecond of physical time. This duration corresponds to $10^{4}$ seconds, namely about three hours of elapsed time over an HPC cluster containing several GPU-based devices Bonaccorso et al. (2022).
@@
At this point, it should be observed that, in one million time steps, the LB populations complete a hundred longitudinal tours along the domain, which sounds very satisfactory. However, the fluid moves much slower than the populations; at a speed of $u=10$ mm/s, in one millisecond, the fluid covers a distance of just ten microns. Hence, the fluid speed must be artificially accelerated by some three orders of magnitude for the fluid to make at least one tour around the device (1 cm long). This is common practice in many simulations, not just LB. The idea is to artificially increase the fluid speed while keeping the relevant dimensionless numbers, in our case the Capillary number $Ca$ and the near-contact number $\mathcal{N}$, unchanged. This can be done by boosting the fluid speed, $u$, the surface tension $\sigma$ and the strength of the near-contact potential $E_{nc}$ by the same factor, so as to keep the ratios $u/\sigma$ and $\alpha/\sigma$ unchanged.
@@ VI.1 Towards quantum nanofluidics
The above scaling derives from the assumption that the fluid molecules in contact with solid walls do not exhibit any net motion, the so called no-slip condition, because the remain trapped in local corrugations of the solid wall. This assumption is no longer valid whenever the size of the channel becomes comparable with the molecular mean free path and, more generally, whenever the fluid-solid molecular interactions cannot be described in terms of simple mechanical collisions.
@@
However, in recent years, it has been argued and experimentally shown that unanticipated quantum-electro-mechanical phenomena can also lead to spectacular reduction of frictional losses Coquinot et al. (2023); Lizée et al. (2023). For instance, it has been shown that slip flow in carbon nanotubes is largely underestimated by MD simulations, indicating that standard force-field procedures fall short of describing the actual physics of the water-graphene interactions, pointing to a crucial role of electronic degrees of freedom in the solid Kavokine et al. (2022). Interestingly, ab-initio MD, besides being even more unpractical computing-wise, would also fall short of capturing the basics of quantum friction because the electrons in the solid are non-adiabatically coupled to charge fluctuations in the liquid. Therefore, it would be extremely interesting to enrich the LB formalism in such a way as to include quantum interfacial interactions. This is indeed possible by exploiting recent results obtained from quantum non-equilibrium Keldysh analysis of interfacial transport in nanofluids Coquinot et al. (2023). The Keldysh formalism accounts for non-equilibrium quantum transport phenomena, and in this respect, it necessarily involves very elaborate analytical calculations based on advanced and retarded Green functions. The final outcome, though, comes in the form of Langevin-like friction terms which could be easily incorporated within an LB-extended framework. For the sake of concreteness let us refer to the case of a charged ionic liquid, in which charge fluctuations couple to phononic and electronic degrees of freedom in the solid walls. The Keldysh analysis shows that, under suitable geometrical and physical conditions, such interfacial quantum coupling can lead to a reduction of the wall friction, because part of the momentum imparted to solid electrons is recycled back to the liquid, an effect known as ”negative quantum friction” (see Fig.22).
@@
Figure 22: Quantum friction scenarios. (a)-(c) A nanofluidic device exhibiting reduced friction through two different electronic current-drive mechanisms: charged fluctuations in liquid water imparting momentum to the electrons in the solid through Coulomb interactions across the liquid-solid interface (b); electrons in the solid driven by phonons, which are excited by water molecule collisions with the solid molecules (c). (d) A sketch of the electron feedback to the liquid. (d) The electron-phonon friction coefficient as a function of the electron density in the solid for four different scenarios: 2D electron gas (2DEG) coupled to water, 2DEG and phonons coupled to water, Graphene coupled to water, graphene with phonons coupled to water. (f) The overall friction coefficient in graphene as compared to classical friction as a function of the electron density. Copyright granted under the https://creativecommons.org/licenses/by/4.0/ CC 4.0 International license. No changes were made from the original figure from Ref.Coquinot et al. (2023).
@@
Clearly, the inclusion of ”quantum friction” alone is not enough, since, based on the fluctuation-dissipation theorem, quantum fluctuations must be accounted for as well. This is a genuinely new theoretical and computational challenge to LB, since, depending on their energy, the quantum fluctuations may last longer then the LB timestep so that, at variance with standard versions of the fluctuating LB formalism Adhikari et al. (2005); Dünweg et al. (2007), the correlator of the fluctuating force extends beyond a single LB time-step.
@@ VI.2 Machine learning
Although ML is essentially a data-driven approach, it can be enriched and reinforced with physical foreknowledge of the problem under investigation. This concept is crucial, for example, in Physics-Informed Machine Learning (PIML) techniques Karniadakis et al. (2021), in which data and mathematical models (such as partial differential or integral equations) are integrated and then implemented through neural networks or kernel-based regression networks Wang et al. (2022); Cai et al. (2021). Thus, ML can be used not only as a heuristic tool driven by the statistical inference of large datasets but also as a system that embeds, in the learning process, the physical laws governing a specific data set. In this framework, the artificial neural networks in the ML approach could act as a class of physical-informed models satisfying universal approximation theorems while preserving the physical ingredients and, concurrently, enhancing the overall computational performance.
@@ VI.2.1 Learning collision operators for ideal fluids
Collision operators for LB simulations of simple fluids can be derived analytically, based on the constraints imposed by the mass-momentum-energy conservation laws. Formally, the mapping from the actual distribution $f$ to the corresponding local equilibrium $f^{eq}$ for the standard single-time BGK collision operator can be written in mode-coupling form as follows:
@@ VI.2.2 Learning generalized collision operators for non-ideal fluids
An intriguing perspective is that offered by the physics-informed neural networks (PINNs). In Ref. Lou et al. (2021), for example, they are used to learn the BGK collision operator of various benchmark flows, such as continuous Taylor-Green flow and rarefied micro-Couette flow. Following a similar footprint, Corbetta et al. Corbetta et al. (2023) have shown that, a neural network architecture combined with suitable physical properties (such as conservation laws and symmetry), allow for a precise reconstruction of collisional operators of lattice Boltzmann and accurate dynamics of standard fluids (such as Taylor-Green vortex and lid-driven cavity flows).
@@
It is known that this search is greatly accelerated by augmenting the loss function with physically inspired constraints. In the case of standard lattice Boltzmann, these are the mass-momentum and momentum-flux constraints. In compact four-dimensional notation, they read as follows:
@@
Further on, with near-contact interactions, the Korteweg tensor must be augmented with the near-contact contributions, as explicitly derived in Montessori et al. (2019a). This procedure requires knowledge of the near-contact interactions augmented Korteweg tensor, hence it is expected to recover existing local equilibria or equivalent forcing terms currently used in LB simulations. However, it cannot be ruled out that new generalized equilibria might emerge from the ML procedure discussed above. This would mark a decisive contribution of machine learning to LB simulations of complex soft matter flows. Although we are not aware of any detailed work in this direction, this is a decidedly interesting topic for future research in the field.

### v2

@@ V.6 Peta-scale simulations of deep-sea sponges
We finally discuss a recent application where the LB method is adopted to study the fluid dynamics of the deep-sea glass sponge Euplectella aspergillum Falcucci et al. (2021), a soft living organism that, besides representing a fascinating example of life under extreme conditions (at depth of $100$-$1000$ meters in the Pacific Ocean and Antarctic area with no ambient sunlight), displays intriguing structural properties due to the trabecular arrangement of its skeletal system Monn et al. (2015); Fernandes et al. (2020).
@@
The interest of this study lies in the fact that, despite the use of a single-component interface-free LB (thus its easiest realization), the detailed three-dimensional reconstruction of the skeleton of the sponge and the computation of the fluid flow in which this organism is immersed need to leverage two of the main advantages of the method, namely the excellent scalability on parallel architectures and the capability of handling complex geometries. Indeed, to reproduce the sponge’s living conditions, the simulations feature over fifty billion grid points spanning approximately four spatial decades, while a complex mesh structure is used to simulate the skeleton.
@@
In Fig.23, we show the computational model and the hydrodynamic flow of the sponge at $Re\simeq 2000$, where $Re=uD/\nu$, being $u$ the water velocity, $\nu$ is its viscosity and $D$ is the sponge diameter at the top section. The sponge structure comprises an anchoring bulb, a section connecting the bulb to the body, the main body and the final sieve plate located at the top (see Fig.23a,b). The body, in particular, is simulated using two intersecting patterns of lattices, the first composed of filaments orthogonally crossing each other and the second made by smaller ligaments placed at $45^{\circ}$ with respect to the first one. The resulting cylindrical structure finally bends into a cone connecting with the anchoring bulb. Also, no-slip conditions are imposed at the internal boundaries of the sponge and at the seafloor.
@@
The simulations show, for example, that a considerable decrease in the flow speed occurs inside the sponge’s body (where vortices are formed) and downstream of the organism, while intermittent fluid patterns emerge several diameters further away (see Fig.23c). These features are supposed to play a role in feeding through a selection mechanism where nutrients are absorbed in the sponge chambers and inorganic particles are discarded. The quiescent region downstream, in particular, is also expected to reduce the hydrodynamic load, an effect quantified in terms of the drag coefficient $C_{D}=2\bar{F}_{drag}/A\rho_{inlet}u^{2}_{inlet}$, where $F_{drag}$ is total drag force, averaged at the steady state and along the flow direction, $\rho_{inlet}$ and $u_{inlet}$ are density and speed at the domain inlet and $A$ is the area of the transverse section of the model, perpendicular to the fluid flow. Indeed, in agreement with previous studies, a considerable drag reduction occurs at $Re>500$ for porous cylindrical models, while the presence of ridges is found to mildly mitigate this effect at large $Re$. More recent simulations have also demonstrated the capability of describing a further aspect of the fluid dynamics inside the body cavity, i.e. the existence of an organized vertical flow toward the oscular aperture, with values of flow speed $\sim 0.83$ cm/s Falcucci et al. (2024), in agreement with experimental results on living Hexatinellida, reporting values around $1$ cm/s Leys et al. (2011). Finally, note that the presence of the sponge’s tissue has been neglected in these simulations although it could impact the structure of the fluid flow, because the tissue may form a barrier with low permeability over the sponge wall Leys et al. (2022).
Figure 23: (a) Drawing of a deep-sea sponge E. aspergillum. (b) A three-dimensional computational model of the organism showing increasing complexity. From left to right: solid cylinder, solid cylinder with helical ridges, porous cylindrical lattice, porous cylindrical lattice with helical ridges, and a complete model. (c) Simulation of a deep-sea sponge immersed in a hydrodynamic flow at $Re\simeq 2000$, where contours of helicity ${\cal H}$ (defined as ${\cal H}={\bf u}\cdot{\bm{\omega}}$, where ${\bm{\omega}}$ is the vorticity) and streaklines of the flow are shown. (d) Drag coefficients at various $Re$ for the simplified models shown in (b) (red circles for solid cylinder, blue squares for solid cylinder with ridges, green diamonds for porous cylinder and orange pentagons for a porous cylinder with ridges) and compared to the following values from literature for cylinders: Ref.Kawamura et al. (1986) pink triangles, Ref.Hanchi et al. (1999) violet diamonds, Ref.Sahin and Owens (2004) red triangles, Ref.N. et al. (2005) empty purple squares, Ref.Henderson (1995) empty light blue pentagons, Ref.Posdziech and Grundmann (2001) empty gray circles. The figure is reproduced with permission from Ref.Falcucci et al. (2021).
V.7 Open challenges.
Despite the fundamental role of the simulations (and specifically LB methods) in the understanding of the properties of these systems, a number of theoretical problems remain open.
In soft flowing crystals and soft granular media (including hierarchical emulsions), for example, a ”comprehensive” multiscale simulation fully covering the spectrum of six orders of magnitude in space and twice as many in time is presently still unfeasible. The reason of such a challenge is that it would require a combination of innovative GPU-based HPC techniques and grid refinement strategies to deal with the necessary computational resources and the increasing algorithmic complexity plus, very likely, new coarse-grained models to improve existing multi-scale theories. Such simulations would be highly desirable for establishing, for example, the extent to which microscopic scale effects (such as interface fluctuations, electrostatic forces, and steric interactions) condition morphology and stability of an emulsion under flow. This is particularly important in systems with high droplet packing fractions (typically higher than the close-packing limit of hard spheres), where interfacial effects often dominate the physics and where refined numerical techniques would be crucial to cope with inevitable numerical instabilities. In these systems, a 3D rheological characterization pinpointing, for example, the role of topological transitions, is currently missing, especially when a highly-packed emulsion crosses narrow gaps, inducing dramatic morphological deformations. These studies, besides being relevant for the design of new soft materials, may also help to describe complex biological processes, such as the collective migration of circulating tumor cells in microchannels Au et al. (2016) or the coordinated movement of hundreds of epithelial cells observed, for example, in wound healing Poujade et al. (2007) and gastrulation Serra et al. (2023).
Intriguing perspectives may also be envisaged for confined dense emulsions subject to thermal flows. For example, it would be interesting to understand the role of heat fluctuations when moving from two to three-dimensional systems, as well as when the number of convective rolls in the Rayleigh-Bénard cell increases. A further open problem deals with the effect of thermal convection in yield stress materials, where the volume fraction of the droplets is very large and the emulsion exhibits non-Newtonian features Pelusi et al. (2024). In this respect, a detailed study of the interplay between the microscopic constituents of these systems and the non-linear rheology is still missing.
In the context of active matter, the study of the physics of self-propelled fluid droplets surely represents a promising field of research. As previously discussed, these systems have been shown to capture a number of features typical of cell motility, such as swimming and crawling. In this respect, a more realistic model than that presented in section V.5 would include a nucleus, whose deformability is known to be a limiting factor for a successful crossing through a micro pore. A minimal mechanistic model could be built using a double emulsion, where the intermediate layer would contain an active liquid crystal while the encapsulated droplet would represent the nucleus. Replacing the solid pillars of the pore with soft ones could be a viable route to adapt this system to the study of more complex biological processes, such as the extravasation in which cancer cells breach the soft barrier of endothelial cells to infiltrate tissues Di Russo et al. (2024).
LB simulations have been also successfully used in the study of biological systems, such as deep-sea glass sponges, for which in vivo experimentation is basically unfeasible. Starting from the results discussed in section V.6, future works may consider the effect of the soft tissue covering the outside surface of the sponge. This condition could be partially reproduced by modeling the sponge as a porous medium with an effective resistance to the flow or through the use of phenomenological boundary conditions capturing the complexity of the combined structure (skeleton plus tissue). However, additional experiments and empirical data are needed to test more complex models than that presented in Refs.Falcucci et al. (2021, 2024).
V.8 From applications to engineering design
@@
The set of selected applications presented in the previous sections is meant to convey the idea of the flexibility and computational efficiency of the LB methods for soft-flowing matter across different regimes of motion. Such flexibility and efficiency are expected to find profitable use in various sectors of industrial and materials design, such as food science, chemical-pharmaceutical (drug design and delivery), tissue engineering and many others. While the details of each and every specific application must necessarily be worked out on a case-by-case basis, in the following we wish to convey the flavor of the potential of computational design in the above sectors based on the LB methods discussed in this review.
@@
For the sake of concreteness, let us refer to a soft material, say a dense emulsion flowing in a box device of height $H=0.1$, width $W=1$ and length $L=10$, all in millimeters. With a lattice spacing $\Delta x=10^{-3}$ (1 micron), the LB simulation consists of $G=10^{2+3+4}=10^{9}$ lattice cells. Such volume contains about a thousand cylindrical droplets of diameter $D=0.1$ mm (100 microns) and height $H$. At a processing speed of $100$ GLUPS, i.e. a hundred billion lattice sites per second, the LB code updates the full box domain a hundred times in a single second of computer time. With a timestep of $\Delta t=10^{-9}$ (1 ns), one million timesteps cover $1$ millisecond of physical time. This duration corresponds to $10^{4}$ seconds, namely about three hours of elapsed time over an HPC cluster containing several GPU-based devices Bonaccorso et al. (2022).
@@
At this point it should be observed that, in one million time steps, the LB populations complete a hundred longitudinal tours along the domain, which sounds very satisfactory. However, the fluid moves much slower than the populations; at a speed of $u=10$ mm/s, in one millisecond, the fluid covers a distance of just ten microns. Hence, the fluid speed must be artificially accelerated by some three orders of magnitude for the fluid to make at least one tour around the device (1 cm long). This is common practice in many simulations, not just LB. The idea is to artificially increase the fluid speed while keeping the relevant dimensionless numbers, in our case the Capillary number $Ca$ and the near-contact number $\mathcal{N}$, unchanged. This can be done by boosting the fluid speed $u$, the surface tension $\sigma$ and the strength of the near-contact potential $E_{nc}$ by the same factor, so as to keep the ratios $u/\sigma$ and $\alpha/\sigma$ unchanged.
@@ VI.1 Towards quantum nanofluidics
The above scaling derives from the assumption that the fluid molecules in contact with solid walls do not exhibit any net motion (the so called no-slip condition), because they remain trapped in local corrugations of the solid wall. This assumption is no longer valid whenever the size of the channel becomes comparable with the molecular mean free path and, more generally, whenever the fluid-solid molecular interactions cannot be described in terms of simple mechanical collisions.
@@
However, in recent years, it has been argued and experimentally shown that unanticipated quantum-electro-mechanical phenomena can also lead to spectacular reduction of frictional losses Coquinot et al. (2023); Lizée et al. (2023). For instance, it has been shown that slip flow in carbon nanotubes is largely underestimated by MD simulations, indicating that standard force-field procedures fall short of describing the actual physics of the water-graphene interactions, pointing to a crucial role of electronic degrees of freedom in the solid Kavokine et al. (2022). Interestingly, ab-initio MD, besides being even more unpractical computing-wise, would also fall short of capturing the basics of quantum friction because the electrons in the solid are non-adiabatically coupled to charge fluctuations in the liquid. Therefore, it would be extremely interesting to enrich the LB formalism in such a way as to include quantum interfacial interactions. This is indeed possible by exploiting recent results obtained from quantum non-equilibrium Keldysh analysis of interfacial transport in nanofluids Coquinot et al. (2023). The Keldysh formalism accounts for non-equilibrium quantum transport phenomena and, in this respect, it necessarily involves very elaborate analytical calculations based on advanced and retarded Green functions. The final outcome, though, comes in the form of Langevin-like friction terms which could be easily incorporated within an LB-extended framework. For the sake of concreteness let us refer to the case of a charged ionic liquid, in which charge fluctuations couple to phononic and electronic degrees of freedom in the solid walls. The Keldysh analysis shows that, under suitable geometrical and physical conditions, such interfacial quantum coupling can lead to a reduction of the wall friction, because part of the momentum imparted to solid electrons is recycled back to the liquid, an effect known as ”negative quantum friction” (see Fig.24).
@@
Figure 24: Quantum friction scenarios. (a)-(c) A nanofluidic device exhibiting reduced friction through two different electronic current-drive mechanisms: charged fluctuations in liquid water imparting momentum to the electrons in the solid through Coulomb interactions across the liquid-solid interface (b); electrons in the solid driven by phonons, which are excited by water molecule collisions with the solid molecules (c). (d) A sketch of the electron feedback to the liquid. (d) The electron-phonon friction coefficient as a function of the electron density in the solid for four different scenarios: 2D electron gas (2DEG) coupled to water, 2DEG and phonons coupled to water, Graphene coupled to water, graphene with phonons coupled to water. (f) The overall friction coefficient in graphene as compared to classical friction as a function of the electron density. Copyright granted under the CC 4.0 International license. No changes were made from the original figure from Ref.Coquinot et al. (2023).
@@
Clearly, the inclusion of ”quantum friction” alone is not enough since, based on the fluctuation-dissipation theorem, quantum fluctuations must be accounted for as well. This is a genuinely new theoretical and computational challenge to LB since, depending on their energy, the quantum fluctuations may last longer then the LB timestep so that, at variance with standard versions of the fluctuating LB formalism Adhikari et al. (2005); Dünweg et al. (2007), the correlator of the fluctuating force extends beyond a single LB time-step.
@@ VI.2 Machine learning
Although ML is essentially a data-driven approach, it can be enriched and reinforced with physical foreknowledge of the problem under investigation. This concept is crucial, for example, in Physics-Informed Machine Learning techniques Karniadakis et al. (2021), in which data and mathematical models (such as partial differential or integral equations) are integrated and then implemented through neural networks or kernel-based regression networks Wang et al. (2022); Cai et al. (2021). Thus, ML can be used not only as a heuristic tool driven by the statistical inference of large datasets but also as a system that embeds, in the learning process, the physical laws governing a specific data set. In this framework, the artificial neural networks in the ML approach could act as a class of physical-informed models satisfying universal approximation theorems while preserving the physical ingredients and, concurrently, enhancing the overall computational performance.
@@ VI.2.1 Learning collision operators for ideal fluids
Collision operators for LB simulations of simple fluids can be derived analytically, based on the constraints imposed by the mass-momentum-energy conservation laws. Formally, the mapping from the actual distribution $f$ to the corresponding local equilibrium $f^{eq}$ for the standard single-time BGK collision operator can be written in mode-coupling form as follows
@@ VI.2.2 Learning generalized collision operators for non-ideal fluids
An intriguing perspective is that offered by the physics-informed neural networks (PINNs). In Ref. Lou et al. (2021), for example, they are used to learn the BGK collision operator of various benchmark flows, such as continuous Taylor-Green flow and rarefied micro-Couette flow. Following a similar footprint, Corbetta et al. Corbetta et al. (2023) have shown that a neural network architecture combined with suitable physical properties (such as conservation laws and symmetry) allow for a precise reconstruction of collisional operators of the LB and accurate dynamics of standard fluids (such as Taylor-Green vortex and lid-driven cavity flows).
@@
It is known that this search is greatly accelerated by augmenting the loss function with physically inspired constraints. In the case of a standard LB, these are the mass-momentum and momentum-flux constraints. In compact four-dimensional notation, they read as follows:
@@
Further on, in the presence of near-contact interactions, the Korteweg tensor must be augmented with these contributions, as explicitly derived in Montessori et al. (2019a). This procedure requires knowledge of the augmented Korteweg tensor, hence it is expected to recover existing local equilibria or equivalent forcing terms currently used in LB simulations. However, it cannot be ruled out that new generalized equilibria might emerge from the ML procedure discussed above. This would mark a decisive contribution of machine learning to LB simulations of complex soft matter flows. Although we are not aware of any detailed work in this direction, this is a decidedly interesting topic for future research in the field.